OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is h = -16t2 + 672t. After how many seconds does it reach its maximum height?

OpenStudy (anonymous):

A. 21s B. 31.5s C. 42s D. 10s

OpenStudy (yrelhan4):

You know how to find maxima using differentiation?

OpenStudy (campbell_st):

do you know how to factor the quadratic of find the line of symmetry

OpenStudy (anonymous):

No, im not sure how to do his problem at all. Ive never been good with word problems.

OpenStudy (anonymous):

If you show me how to find the answer to this practice problem, ill be bale to figure out the answer to the actual problem i need

OpenStudy (campbell_st):

well the line of symmetry for a parabola \[y = ax^2 + bx + c\] is given by \[x = \frac{-b}{2a}\] the maximum height occurs on the line of symmetry so in your question you have to find \[t = \frac{-b}{2a} \] where b = 672 and a = -16 this will give the time when the maximum height occurs